# DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering

### v1

@@ Appendix C Additional results
BERTScore relies on token-level semantic similarity, which rewards shallow lexical overlap rather than actual factual accuracy. For example, in cases where minor differences in wording (e.g., “The Treaty of Versailles was signed in 1919.” versus “The Treaty of Versailles ended in 1919.”) lead to opposing factual claims, BERTScore still scores the response high due to its emphasis on matching tokens (e.g., “signed” versus “ended”). Additionally, verbosity bias and threshold instability—where a default threshold (threshold = 0.5) is arbitrarily set—further inflate its raw accuracy. However, when comparing raw accuracy with instance-level agreement metrics like Cohen’s kappa, which adjusts for class imbalance and penalizes asymmetric errors, the limitations of BERTScore become apparent.
LLMs Tasks Evaluators
EM BS HM DeepSeek Llama GPT Mixtral Mistral
DeepSeek AmbigQA 56.3 80.0 84.3 86.3 73.7 75.0 62.3 93.3
FreshQA 31.3 88.0 84.3 84.7 82.7 75.3 58.0 82.3
HotpotQA 38.6 78.4 57.7 58.0 51.0 51.0 52.7 57.7
NQ-Open 35.0 78.3 60.3 64.7 63.7 61.3 55.3 68.3
TriviaQA 77.3 90.7 94.3 90.7 94.0 91.7 81.7 89.7
Llama AmbigQA 42.3 63.0 67.0 64.0 65.3 64.7 63.0 66.0
FreshQA 25.6 81.3 77.7 81.3 78.3 72.7 71.0 62.3
HotpotQA 34.3 67.7 56.3 56.7 58.3 54.0 50.7 52.7
NQ-Open 31.7 61.7 66.3 62.3 62.7 60.0 59.0 66.7
TriviaQA 74.3 94.0 94.7 88.0 90.3 90.0 88.7 84.7
GPT AmbigQA 49.7 78.0 71.7 70.3 70.0 68.0 65.7 71.0
FreshQA 24.6 89.3 70.7 58.0 51.7 78.7 83.0 83.3
HotpotQA 33.7 80.0 54.0 50.3 53.0 52.7 51.7 54.0
NQ-Open 36.3 74.0 65.3 65.3 62.7 59.0 59.0 67.0
TriviaQA 74.3 95.3 93.0 90.0 89.3 90.7 89.7 86.3
Mixtral AmbigQA 37.7 70.3 61.7 58.7 57.3 62.0 59.3 61.7
FreshQA 18.6 89.7 86.0 72.3 67.0 87.0 85.0 77.7
HotpotQA 25.0 69.7 47.0 46.3 45.3 45.7 44.7 46.0
NQ-Open 23.7 63.7 56.7 54.0 52.7 47.7 52.3 59.7
TriviaQA 64.7 91.3 90.7 83.7 86.3 89.7 86.0 85.3
Mistral AmbigQA 31.0 61.7 49.7 47.7 46.3 47.7 46.3 53.3
FreshQA 15.6 80.0 81.7 60.7 59.0 83.7 84.0 86.0
HotpotQA 23.7 64.7 40.0 39.3 39.0 38.0 37.0 39.0
NQ-Open 22.7 60.0 46.0 41.3 40.0 43.3 41.3 50.0
TriviaQA 62.0 94.3 83.7 78.0 81.3 81.0 79.7 85.0
Table 6: Raw performance of candidate LLMs across free-form QA tasks evaluated through various methods. HM represents Human Majority and BS denotes BERTScore.
C.1 Impact of arbitration on dispute resolution
Figure 6 illustrates the impact of arbitration on resolving disagreements between primary judges. Arbitration, facilitated by GPT-3.5 as the tiebreaker, consistently improves performance across all tasks, particularly in FreshQA and TriviaQA, where Macro F1 increases by up to 21.5 points. In contrast, tasks like AmbigQA and HotpotQA, where primary judges initially exhibit stronger agreement, show smaller but still meaningful improvements. This highlights the critical role of arbitration in enhancing agreement and achieving closer alignment with ground truth, especially in cases of significant disagreement among primary judges.
Notably, evaluations of DeepSeek-v3 exhibit higher disagreement between Llama-3.1-70B and Mistral-7B, particularly in FreshQA (28.3%) and AmbigQA (25.7%). From our analysis, we did not find strong evidence explaining why DeepSeek-v3 leads to higher disagreement between the primary judges.
Figure 6: Impact of arbitration on disagreements between primary judges. Note that we used Llama-3.1-70B and Mistra 7B as primary judges. GPT-3.5-turbo is only utilized when disagreements are found. The models given in the figure are candidate LLMs which generate outputs for the given tasks and are then evaluated through DAFE.
We observed substantial enhancements in Cohen’s Kappa scores across several tasks. For instance, as illustrated in Figure 7, in the AmbigQA Cohen’s Kappa increased from 0.881 to 0.911 for Llama. Similarly, in the same task, Cohen’s Kappa from 0.467 to 0.773 for candidate DeepSeek. These improvements demonstrate that the arbitration mechanism effectively enhances the reliability and consistency of evaluations, particularly in complex and ambiguous tasks where primary judges are more likely to disagree.
Some Cohen’s Kappa scores remain relatively low, particularly in FreshQA and DeepSeek-evaluated outputs. This is partially explained by the Kappa Paradox, where high agreement on extreme cases (e.g., clear correct/incorrect responses) and unbalanced class distributions can artificially lower the Kappa scores. In such cases, even when evaluators mostly agree, Cohen’s Kappa can appear lower than expected. Despite this, the arbitration process effectively mitigates inconsistencies, especially in tasks involving evolving knowledge and nuanced interpretations, such as FreshQA.
Figure 7: Comparison of Cohen’s kappa scores before and after arbitration (GPT-3.5-turbo as arbitrator). The performance is illustrated across candidate LLMs and tasks.
C.2 Cost analysis
Human evaluation is the gold standard for assessing LLM-generated responses, but it is expensive and time-consuming. In our setup, we employed three human annotators who volunteered their efforts. However, if these annotators were compensated based on standard annotation rates, the cost of evaluating such outputs would be significantly higher. On the other hand, GPT-3.5-turbo, acting as an arbitrator in DAFE, incurs a cost that depends on the number of arbitration cases. In our evaluations, GPT-3.5-turbo was invoked 1,318 times, with an estimated total cost of \$0.59, which increases to \$5.40 if a 2048 max token setting is used (see Table 7). Since GPT-3.5 is only invoked when primary judges disagree, this selective arbitration substantially reduces overall evaluation expenses while maintaining high reliability in assessments. Rather than relying on a single model for evaluation, this multi-model arbitration approach enhances trust by mitigating biases and weaknesses inherent in any individual model.
By invoking the arbitrator only when disagreements occur (rather than evaluating all responses), DAFE reduces arbitration usage by 82–94% compared to a majority-voting system. This leads to:
  • Over 90% fewer third-judge inferences, drastically lowering computational demand.
  • Up to 95% cost savings by avoiding redundant model evaluations.
  • Better scalability, making it practical for large-scale deployments
Table 7: Cost-efficiency analysis of DAFE: Summary of disagreement rates and tiebreaker usage across candidate models and tasks
@@ C.2 Cost analysis
Candidate LLMs Tasks Samples Disagreement Rates (%) Tiebreaker Usage
@@
C.3 DeepSeek as the arbitrator
To assess the impact of using DeepSeek as the arbitrator in DAFE, we conducted experiments by replacing GPT-3.5-turbo with DeepSeek. We evaluated this setup using different candidate models across multiple tasks. Specifically, we tested GPT-3.5 on TriviaQA, DeepSeek on NQ-Open, and Llama on FreshQA. The primary judges remained Llama and Mistral, and arbitration was invoked only in cases of disagreement. Our findings indicate that DeepSeek as the arbitrator achieves strong performance, with Macro-F1 scores of 91.23 on TriviaQA, 79.11 on NQ-Open, and 0.914 on FreshQA.
C.4 Evaluating with one strong LLM-as-a-judge
@@
To explore the potential of a more powerful single LLM, we evaluated GPT-3.5-turbo on HotpotQA and TriviaQA using GPT-4o as a judge. With this configuration, GPT-4o as the evaluator achieved a Macro-F1 score of 0.946 on HotpotQA, demonstrating its exceptional capability. However, the same GPT-4o judge achieved only 0.784 on TriviaQA, which falls short of DAFE’s performance of 0.887. This shows that even the most advanced models show inconsistencies when evaluating free-form QA. This is particularly critical in precision-sensitive domains where minor errors can have outsized consequences.
In such settings, DAFE’s ensemble approach acts as a safeguard. When employing DAFE with GPT-3.5-turbo as the arbitrator, we achieved an even higher Macro-F1 of 0.984 on HotpotQA, surpassing the performance of a single GPT-4o. Interestingly, when we experimented with DeepSeek as the arbitrator in DAFE, performance remained strong at 0.963 Macro-F1, indicating that DAFE’s benefits are not solely tied to a specific arbitrator model.
C.5 Majority voting-based evaluation
We conducted additional experiments utilizing a traditional majority voting approach for evaluating candidate LLM performance. In this setup, we employed three LLM judges of equal weight: Llama, GPT-3.5, and Mistral to evaluate candidate models generated response. For every evaluation instance, each judge provided an independent binary verdict (True or False). The final decision is determined through a simple majority vote across these three verdicts.
As presented in Table 8, DAFE matches or closely approaches the Macro F1 and Cohen’s Kappa scores of the three-judge majority across almost all tasks and candidate LLMs. For example, on HotpotQA, evaluating candidate Llama with DAFE achieves a Macro F1 of 97.6% (compared to 97.6% for majority voting) and a Cohen’s Kappa of 0.95, while for GPT-3.5 on AmbigQA, DAFE reaches a Macro F1 of 98.4% (versus 98.3% for majority voting), indicating a negligible performance difference. Even in high-disagreement tasks like TriviaQA, where the primary judges (e.g., Mistral) disagree 20.3% of the time, DAFE retains strong alignment (with a Macro F1 of 92.7 compared to 93.5 for majority voting). Minor deviations, such as the one observed for candidate Mixtral on TriviaQA (DAFE’s Macro F1 = 0.88 vs. 0.95 for majority voting), reflect rare instances where both the primary judges and the arbitrator make errors, yet these outliers are substantially outweighed by the computational savings offered by selective arbitration.
Table 8: Comparison between Majority Voting (Llama+GPT-3.5+Mistral) and DAFE (GPT-3.5 as arbitrator). For each candidate LLM and task, the table reports Macro F1 and Cohen’s Kappa scores under Majority Voting, the disagreement rate (in %), and the corresponding scores using DAFE.
Candidate LLM Task Majority Voting Disagreement (%) DAFE
@@ C.5 Majority voting-based evaluation
C.6 Impact of prompt variations
@@
C.6.1 Consistency in judgment across multiple trials
LLMs generate random text even at a temperature of 0. To assess whether this affects evaluation consistency, we repeated the same evaluation task five times for 100 Mistral-generated responses for HotpotQA.
  • Verdict stability: GPT produced identical True/False verdicts in 100% of cases. This suggest that its binary decision-making process remains stable even across multiple trials.
@@ C.6.1 Consistency in judgment across multiple trials
C.6.2 Few-shot vs. zero-shot prompting
@@
C.6.3 Explanation requirement: Binary verdict vs. justification-based evaluation
@@
C.6.4 Reason-first vs. verdict-first prompting
@@
C.7 G-Eval: reference-free evaluation of free-form question-answering
Existing LLM-based evaluators such as G-Eval (Liu et al., 2023) are designed for reference-free, subjective tasks (e.g., summarization, dialogue), where evaluation criteria (e.g., coherence, fluency) are inherently ambiguous and scored on Likert scales. These frameworks prioritize qualitative judgments rather than binary factual correctness. In contrast, DAFE is explicitly tailored for reference-dependent, objective evaluation in free-form QA, where answers are either factually correct or incorrect based on alignment with explicit ground-truth references.
To validate this distinction, we tailored G-eval based method to investigate the capability of LLM-as-a-judge in reference-free settings. In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions such as correctness.
The performance of LLM-as-a-judge drastically changes in reference-free settings. Without access to the ground truth references, we observe a stark decline in evaluation capability across all models (see Table 9 and 10 values in blue). This systematic deterioration spans all tasks and model combinations, though its severity varies by context. HotpotQA, with its demands for complex reasoning, exemplifies this challenge most clearly. The substantial gap between reference-based and reference-free evaluation underscores the crucial role of reference answers in reliable assessment.
@@ C.7 G-Eval: reference-free evaluation of free-form question-answering
Table 9: Overall performance (Accuracy) of candidate LLMs across free-form QA tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mood.
@@
EM BERTScore Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B
Llama-3.1-70B AmbigQA 0.744 0.641 0.944 [0.629] 0.922 [0.604] 0.912 [0.669] 0.929 [0.631]
HotpotQA 0.778 0.745 0.939 [0.628] 0.949 [0.574] 0.910 [0.665] 0.916 [0.640]
NQ-Open 0.653 0.718 0.916 [0.606] 0.896 [0.560] 0.907 [0.639] 0.869 [0.622]
TriviaQA 0.612 0.782 0.772 [0.772] 0.717 [0.628] 0.695 [0.678] 0.640 [0.633]
GPT-3.5 AmbigQA 0.792 0.622 0.972 [0.686] 0.949 [0.603] 0.930 [0.596] 0.927 [0.553]
HotpotQA 0.794 0.623 0.977 [0.566] 0.987 [0.521] 0.936 [0.543] 0.966 [0.494]
NQ-Open 0.703 0.606 0.942 [0.671] 0.911 [0.544] 0.911 [0.601] 0.914 [0.536]
TriviaQA 0.646 0.681 0.824 [0.817] 0.700 [0.690] 0.789 [0.760] 0.730 [0.701]
Mixtral-8x7B AmbigQA 0.760 0.666 0.948 [0.704] 0.891 [0.636] 0.955 [0.654] 0.944 [0.622]
HotpotQA 0.761 0.657 0.970 [0.587] 0.966 [0.470] 0.930 [0.582] 0.970 [0.577]
NQ-Open 0.650 0.649 0.939 [0.652] 0.863 [0.517] 0.950 [0.590] 0.908 [0.529]
TriviaQA 0.625 0.695 0.812 [0.800] 0.803 [0.754] 0.838 [0.818] 0.716 [0.725]
Mistral-7B AmbigQA 0.792 0.622 0.947 [0.730] 0.947 [0.627] 0.947 [0.628] 0.930 [0.523]
HotpotQA 0.796 0.673 0.969 [0.649] 0.951 [0.478] 0.947 [0.680] 0.969 [0.578]
NQ-Open 0.726 0.639 0.925 [0.652] 0.919 [0.515] 0.939 [0.597] 0.920 [0.433]
TriviaQA 0.718 0.608 0.879 [0.881] 0.863 [0.840] 0.822 [0.846] 0.735 [0.744]
Table 10: Performance (Macro F1) of various evaluators across candidate LLMs and tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mode.
C.8 DAFE in multi-reference answers
DAFE explicitly accommodates multiple gold reference answers by incorporating all available references into the judge LLM’s prompt during evaluation. For datasets like AmbigQA and TriviaQA, where questions often have multiple valid answers (e.g., synonyms, rephrased answers, or alternative factual representations), DAFE aggregates all reference answers into the judge’s input prompt (e.g., concatenating them as a comma-separated list).

### v2

@@ Appendix E Additional results
BERTScore relies on token-level semantic similarity, which rewards shallow lexical overlap rather than actual factual accuracy. For example, in cases where minor differences in wording (e.g., “The Treaty of Versailles was signed in 1919.” versus “The Treaty of Versailles ended in 1919.”) lead to opposing factual claims, BERTScore still scores the response high due to its emphasis on matching tokens (e.g., “signed” versus “ended”). Additionally, verbosity bias and threshold instability further inflate its raw accuracy (see Table 10). However, when comparing raw accuracy with instance-level agreement metrics like Cohen’s kappa, which adjusts for class imbalance and penalizes asymmetric errors, the limitations of BERTScore become apparent.
E.1 Cost analysis
To assess the efficiency of CLEV, we track the number of times the third judge is invoked, which directly corresponds to disagreement between the two primary judges. As shown in Table 8, across 7,500 evaluation instances, the third judge is required only 1,318 times, representing just 17.6% of the total cases. This implies an 82.4% reduction in third-judge usage compared to a full majority-voting setup, where every instance would involve all three models.
Disagreement rates vary across tasks and models. For example, GPT shows only 5.7% disagreement on HotpotQA, while FreshQA exhibits higher disagreement (up to 44.3%) for some judge combinations. This behavior allows CLEV to scale efficiently: it concentrates computational effort only where model uncertainty exists, minimizing redundant inference. In contrast to fixed-cost evaluation schemes, CLEV offers a cost-efficient alternative that maintains high evaluation quality while significantly reducing compute usage.
@@ E.1 Cost analysis
Candidate LLMs Tasks Samples Disagreement Rates (%) Third Judge Usage
@@
Table 8: Cost-efficiency analysis of CLEV: Summary of disagreement rates and third judge usage across candidate models and tasks
E.2 DeepSeek as the third judge
To assess the impact of using DeepSeek as the third judge in CLEV, we conducted experiments by replacing GPT-3.5-turbo with DeepSeek-R1 (deepseekai2025deepseekr1). We evaluated this setup using candidate GPT-3.5 on TriviaQA, candidate DeepSeek on NQ-Open, and candidate Llama on FreshQA. The primary judges remained Llama and Mistral, and the third judge (i.e., DeepSeek-R1) is invoked only in cases of disagreement. Our findings indicate that DeepSeek, as the third judge, achieves strong performance, with Macro-F1 scores of 91.23 on TriviaQA, 79.11 on NQ-Open, and 0.914 on FreshQA.
E.3 Evaluating with GPT-4o as-a-judge
@@
To explore the potential of a more powerful single LLM, we evaluated GPT-3.5-turbo on HotpotQA and TriviaQA using GPT-4o as a judge. With this configuration, GPT-4o as the evaluator achieved a Macro-F1 score of 0.946 on HotpotQA, demonstrating its exceptional capability. However, the same GPT-4o judge achieved only 0.784 on TriviaQA, which falls short of CLEV’s performance of 0.887. This shows that even the most advanced models show inconsistencies when evaluating free-form QA. This is particularly critical in precision-sensitive domains where minor errors can have outsized consequences.
In such settings, CLEV’s ensemble approach acts as a safeguard. When employing CLEV with GPT-3.5-turbo as the third judge, we achieved an even higher Macro-F1 of 0.984 on HotpotQA, surpassing the performance of a single GPT-4o. Interestingly, when we experimented with DeepSeek as the third judge in CLEV, performance remained strong at 0.963 Macro-F1, indicating that CLEV’s benefits are not solely tied to a specific third judge model.
E.4 Majority voting-based evaluation
We conducted additional experiments utilizing a traditional majority voting approach for evaluating candidate LLMs performance. Given $n$ annotators and a binary classification, the majority label is defined as:
$y_{\text{majority}}=\begin{cases}1&\text{if }\sum_{i=1}^{n}y_{i}>\frac{n}{2},\\ 0&\text{otherwise},\end{cases}$
where $y_{i}$ represents the label assigned by the $i$th annotator.
In this setup, we employed three LLM judges of equal weight: Llama, GPT-3.5, and Mistral to evaluate candidate models’ generated responses. For every evaluation instance, each judge provided an independent binary verdict (True or False). The final decision is determined through a simple majority vote across these three verdicts.
As presented in Table 9, CLEV matches or closely approaches the Macro F1 and Cohen’s Kappa scores of the three-judge majority across almost all tasks and candidate LLMs. For example, on HotpotQA, evaluating candidate Llama with CLEV achieves a Macro F1 of 97.6% (compared to 97.6% for majority voting) and a Cohen’s Kappa of 0.95, while for GPT-3.5 on AmbigQA, CLEV reaches a Macro F1 of 98.4% (versus 98.3% for majority voting), indicating a negligible performance difference. Even in high-disagreement tasks like TriviaQA, where the primary judges (e.g., Mistral) disagree 20.3% of the time, CLEV retains strong alignment (with a Macro F1 of 92.7 compared to 93.5 for majority voting). Minor deviations, such as the one observed for candidate Mixtral on TriviaQA (CLEV’s Macro F1 = 0.88 vs. 0.95 for majority voting), reflect rare instances where both the primary judges and the third judge make errors, yet these outliers are substantially outweighed by the computational savings offered by the selective third judge.
Candidate LLM Task Majority Voting Disagreement (%) CLEV
@@ E.4 Majority voting-based evaluation
Table 9: Comparison between Majority Voting (Llama+GPT-3.5+Mistral) and CLEV (GPT-3.5 as the third judge). For each candidate LLM and task, the table reports Macro F1 and Cohen’s Kappa scores under Majority Voting, the disagreement rate (in %), and the corresponding scores using CLEV. Small differences in performance between the majority voting setup and CLEV are due to randomness (see Section E.5.1).
E.5 Impact of prompt variations
@@
E.5.1 Consistency in judgment across multiple trials
LLMs generate random text even at a temperature of 0. To assess whether this affects evaluation consistency, we repeated the same evaluation task five times for 300 Mistral-generated responses for HotpotQA. In this ablation study, we use GPT-3.5 as a judge.
  • Verdict stability: GPT produced identical True/False verdicts in 98% of cases, with minor variations observed in only 1–2 instances. This suggest that its binary decision-making process remains largely stable even across multiple trials.
@@ E.5.1 Consistency in judgment across multiple trials
E.5.2 Few-shot vs. zero-shot prompting
@@
E.5.3 Explanation requirement: Binary verdict vs. justification-based evaluation
@@
E.5.4 Reason-first vs. verdict-first prompting
@@
E.6 G-Eval: reference-free evaluation of free-form question-answering
Existing LLM-based evaluators such as G-Eval (liuetal2023g) are designed for reference-free, subjective tasks (e.g., summarization, dialogue), where evaluation criteria (e.g., coherence, fluency) are inherently ambiguous and scored on Likert scales. These frameworks prioritize qualitative judgments rather than binary factual correctness. In contrast, CLEV is explicitly tailored for reference-dependent, objective evaluation in free-form QA, where answers are either factually correct or incorrect based on alignment with explicit ground-truth references.
To validate this distinction, we tailored the G-eval (liuetal2023g) to investigate the capability of LLM-as-a-judge in reference-free settings. In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompt the evaluator model as $P=\{x,\bar{y}\}$ along with instructions such as correctness.
The performance of LLM-as-a-judge drastically changes in reference-free settings. Without access to the ground truth references, we observe a stark decline in evaluation capability across all models (see Table 10 and 11 values in blue). This systematic deterioration spans all tasks and model combinations, though its severity varies by context. HotpotQA and NQ-Open, with their demands for complex reasoning, exemplify this challenge most clearly. The substantial gap between reference-based and reference-free evaluation underscores the crucial role of reference answers in reliable assessment.
@@ E.6 G-Eval: reference-free evaluation of free-form question-answering
Table 10: Overall performance (Raw Accuracy) of candidate LLMs across free-form QA tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mood.
@@
EM BERTScore Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B CLEV
Llama-3.1-70B AmbigQA 0.744 0.641 0.944 [0.629] 0.922 [0.604] 0.912 [0.669] 0.929 [0.631] 0.955 [0.637]
HotpotQA 0.778 0.745 0.939 [0.628] 0.949 [0.574] 0.910 [0.665] 0.916 [0.640] 0.976 [0.623]
NQ-Open 0.653 0.718 0.916 [0.606] 0.896 [0.560] 0.907 [0.639] 0.869 [0.622] 0.964 [0.610]
TriviaQA 0.612 0.782 0.772 [0.772] 0.717 [0.628] 0.695 [0.678] 0.640 [0.633] 0.842 [0.747]
GPT-3.5 AmbigQA 0.792 0.622 0.972 [0.686] 0.949 [0.603] 0.930 [0.596] 0.927 [0.553] 0.984 [0.607]
HotpotQA 0.794 0.623 0.977 [0.566] 0.987 [0.521] 0.936 [0.543] 0.966 [0.494] 0.993 [0.522]
NQ-Open 0.703 0.606 0.942 [0.671] 0.911 [0.544] 0.911 [0.601] 0.914 [0.536] 0.978 [0.575]
TriviaQA 0.646 0.681 0.824 [0.817] 0.700 [0.690] 0.789 [0.760] 0.730 [0.701] 0.887 [0.882]
Mixtral-8x7B AmbigQA 0.760 0.666 0.948 [0.704] 0.891 [0.636] 0.955 [0.654] 0.944 [0.622] 0.975 [0.650]
HotpotQA 0.761 0.657 0.970 [0.587] 0.966 [0.470] 0.930 [0.582] 0.970 [0.577] 0.987 [0.536]
NQ-Open 0.650 0.649 0.939 [0.652] 0.863 [0.517] 0.950 [0.590] 0.908 [0.529] 0.956 [0.563]
TriviaQA 0.625 0.695 0.812 [0.800] 0.803 [0.754] 0.838 [0.818] 0.716 [0.725] 0.882 [0.858]
Mistral-7B AmbigQA 0.792 0.622 0.947 [0.730] 0.947 [0.627] 0.947 [0.628] 0.930 [0.523] 0.977 [0.647]
HotpotQA 0.796 0.673 0.969 [0.649] 0.951 [0.478] 0.947 [0.680] 0.969 [0.578] 0.979 [0.673]
NQ-Open 0.726 0.639 0.925 [0.652] 0.919 [0.515] 0.939 [0.597] 0.920 [0.433] 0.976 [0.527]
TriviaQA 0.718 0.608 0.879 [0.881] 0.863 [0.840] 0.822 [0.846] 0.735 [0.744] 0.927 [0.913]
Table 11: Performance (Macro F1) of various evaluators across candidate LLMs and tasks. Values [in blue] represent the reference-free mode.
E.7 CLEV in multi-reference answers
CLEV explicitly accommodates multiple gold reference answers by incorporating all available references into the judge LLM’s prompt during evaluation. For datasets like AmbigQA and TriviaQA, where questions often have multiple valid answers (e.g., synonyms, rephrased answers, or alternative factual representations), CLEV aggregates all reference answers into the judge’s input prompt (e.g., concatenating them as a comma-separated list).